# Rotating axionic AdS4 black hole dressed with a scalar field

### v1

@@ IV Holographic superconductor
In this section, we are interested in exploring a holographic superconductor description of the preceding BH configuration. Given this, we find it convenient to focus on a particular case of this background in which the transformation $u\to r_{h}/r$ makes it feasible to locate the outer horizon at $u=1$ and the asymptotic boundary at $u=0$. This is possible, provided we define
@@
which preserves the genuine nature of the parameter $\alpha$ as a rotation parameter. In this manner, by evaluating the above expressions into the background functions (12)-(19) and taking the limit $\alpha\to 0^{+}$, the angular momentum of the background is turned off and a static AdS4 BH configuration is obtained (see eqs. (20)-(21)). Moreover, in the setting (46), the Hawking temperature (44) adopts the simpler form
@@
in which $F_{\mu\nu}=2\partial_{[\mu}A_{\nu]}$, while $\Psi$ denotes a complex scalar field characterized by mass $m$ and charge $q$, minimally coupled to the gauge field $A_{\mu}$ through the extended covariant derivative $D_{\mu}=\nabla_{\mu}-iA_{\mu}$. Here, we work in the probe limit $q\gg 1$, where the backreaction of the matter contribution (48) onto the background fields can be effectively ignored. Thus, the decoupled system of field equations is given by
@@
which is intended to be solved on the fixed background (6) taking into account eq. (46).
@@ IV.1 Condensation of the scalar field and the electric conductivity
where now the primes $(^{\prime})$ stand for derivatives with respect to $u$. In solving this system of equations, we need to specify appropriate boundary conditions at the horizon ($u=1$) and at the conformal boundary of the background ($u=0$). Concretely, at the horizon, we require that the scalar field is finite $\Psi(1)<\infty$, while the components $A_{t}(1)$ and $A_{\theta}(1)$ vanish for $A_{\mu}$ to have a finite norm everywhere. On the other hand, at the boundary, the fields behave as
@@
From the holographic standpoint, the coefficient $\Psi_{1}$ (respectively $\Psi_{2}$) is identified as the order parameter $\mathcal{O}$ with scaling dimension $\Delta_{1}$ (respectively $\Delta_{2}$). However, to have well-defined normalizable modes at the Ultraviolet (UV) regime (boundary), we must restrict the scalar field mass to be bounded below as $m^{2}\ell^{2}\geq-\frac{9}{4}$, which defines the Breitenlohner-Freedman (BF) bound for a massive scalar field in AdS4 Breitenlohner:1982bm ; Breitenlohner:1982jf . Besides, $\mu$ and $\rho$ are respectively interpreted as the chemical potential and the charge density operator. A similar interpretation is given to the coefficients of $A_{\theta}$, where $\mu_{\theta}$ stands for a one-dimensional potential vector, source of the density of current operator $\mathcal{J_{\theta}}$.
@@
Using the holographic dictionary, by numerically solving the system (51), we can determine the condensation profile of the order parameter as a function of the temperature. For our analysis, we adopt the canonical Ensemble, maintaining a constant charge density $\rho$. In addition, we fix the mass to $m^{2}\ell^{2}=-2$, a value consistent with the BF bound and that gives the simplest asymptotic structure for the scalar field. This choice sets the scaling dimensions $\Delta_{1}=1$ and $\Delta_{2}=2$ for the dual operators $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$, respectively. Fig. 2 depicts the effect of the BH rotation on the condensation curve for $\mathcal{O}_{1}$ (left panel) and $\mathcal{O}_{2}$ (right panel). As shown, increasing the BH rotation (angular momentum) leads to a reduction in the amplitude of the condensation operator. This behavior aligns with findings from previous holographic models that also incorporates rotating BH backgrounds albeit with a different asymptotic nature Lin:2014tza ; Herrera-Mendoza:2024vfj ; Srivastav:2019ixc .
@@
Figure 2: The condensation profiles for the operators $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ as a function of temperature, considering different values of the rotation parameter $\alpha$. The continuous (black) curve corresponds to the static case, while the dashed (violet), dot-dashed (blue), and dotted (red) correspond to situations with non-zero rotation.
In addition to the above, we present the electric conductivity $\sigma_{\tiny{\mbox{DC}}}$, where the idea is to analyze the linear response of the system to a small perturbation of external sources. Concretely, via fluctuations of the vector potentials $A_{\theta}$ and $A_{\varphi}$, denoted as $\delta A_{\theta}$ and $\delta A_{\varphi}$ on the boundary, and the retarded Green’s function. For this, we need to employ the fluctuations (see Refs. Melnikov:2012tb ; Hartnoll:2008vx ):
$\delta A_{\theta,\varphi}=A_{\theta,\varphi}(r)e^{-i\omega\,t},\quad\delta\phi=\delta\psi=0,$ (54)
where $A_{\theta,\varphi}$ has been introduced to jointly represent the components $A_{\theta}$ and $A_{\varphi}$, while that $\omega$ corresponds to the frequency. Considering the complex fluctuation $A=A_{\theta}d\theta+iA_{\varphi}d\varphi$ on the field equations $\nabla_{\mu}F^{\mu\nu}=0$, the equations of motion read
$A{{}^{\prime\prime}}_{\theta,\varphi}+\frac{1}{2}\left(\frac{f{{}^{\prime}}}{f}+\frac{h{{}^{\prime}}}{h}\right)A{{}^{\prime}}_{\theta,\varphi}+\frac{\omega^{2}}{fh}A_{\theta,\varphi}=0,$ (55)
@@
Then, to extract $\sigma_{\tiny{\mbox{DC}}}$, which can be calculated through linear response techniques Hartnoll:2008vx ; Hartnoll:2008kx , given an external oscillating electric field $E_{i}(\omega)$ generating a corresponding current $J_{i}(\omega)$, with $i=\{\theta,\varphi\}$, the electrical conductivity is defined as the ratio
$\sigma_{\tiny{\mbox{DC}}}(\omega)=\frac{J_{\theta,\varphi}(\omega)}{E_{\theta,\varphi}(\omega)},$ (56)
where we can compute the electric conductivity through the Kubo formula as a two-point function Baggioli:2016rdj :
@@
$\sigma_{\tiny{\mbox{DC}}}(\omega)=-\frac{i}{\omega}G^{R}_{J_{\theta,\varphi}J_{\theta,\varphi}}=\frac{A{{}^{\prime}}_{\theta,\varphi}}{i\omega A_{\theta,\varphi}}.$ (57)
@@
Figure 3: The $\sigma_{\tiny{\mbox{DC}}}$ curve for the real (left panel) and imaginary (right panel) parts of the electrical conductivity for different values of the rotation parameter $\alpha$. The continuous (black) curve corresponds to the static case, while the other values correspond to situations with non-zero angular momentum.
Figure 3 shows an example of quantum holographic of electronic wave functions, which appears when rotation parameter $\alpha$ increases. This behavior shows us a macroscopic superconducting condensate phase Santos:2024cwf ; Hovhannisyan:2022gkp . We note that the situation $\omega=0$ represents a divergence characterizing a pole in the imaginary part of the conductivity Baggioli:2016rdj . The breaking of symmetry leads to the non-conservation of momentum. This way, the charges dissipate their momentum, and the current will not persist forever. Thus, a physical consequence of these results for real materials is the appearance of impurities, lattice, and disorder, due to the breakdown of momentum conservation Amoretti:2019cef ; Ammon:2019wci ; Andrade:2017lsc ; Tanhayi:2016uui .
@@ V Conclusions and discussions
In the present work, we construct an axionically charged spinning black hole in four dimensions with a non-minimally coupled scalar field. The above is possible via the inclusion of a suitable potential as well as a structural function coupling to the axionic field. It is interesting to note that, although the metric function $f$ and the rotating function $N^{\theta}$ enjoy logarithmic behavior, this configuration is asymptotically AdS.
As can be seen in Fig. 1 (left panel), this configuration is characterized by two parameters ($\alpha$ and $\eta$) and only one integration constant $B$, allowing us to obtain one or even two horizons. Furthermore, we find that the metric function $f$ and the scalar field $\phi$ do not vanish for $\alpha\rightarrow 0^{+}$ and $\eta\neq 0$. This allows us to obtain a new static hairy BH configuration with one horizon (see Fig. 1 (right panel)).
Additionally, due to the inclusion of the scalar and axionic field, this model allows us the emergence of new thermodynamic parameters through the Euclidean action (30), these are the mass $\mathcal{M}$, Hawking temperature $T$, angular momentum $\mathcal{J}$, angular velocity $\Omega$, as well as the axionic potential $\psi_{a}$, and axionic charge $\omega_{a}$, satisfying first law (45) and the four-dimensional Smarr relation Smarr
@@
In exploring the holographic superconductor description of these spinning BHs, we employed a numerical approach to investigate the condensation of the order parameter as a function of temperature and the rotation parameter (see Fig. 2). Notably, the logarithmic falloff of the blackening function does not alter the typical profile of the condensate. Nevertheless, our results reveal that the BH rotation impacts the order parameter. Specifically, we found that increasing the rotation diminishes the amplitude of the condensation curve, suggesting that, for the chosen parameter values, the presence of rotation suppresses the superconducting state in our holographic system. This effect becomes relevant only at sufficiently high values of the rotation parameter. Together with the aforementioned, the electric conductivity as a function of the frequency $\omega$ is calculated numerically from the spinning configuration through the response of the system to small external sources. This is followed by the calculation of the Green’s functions for the electric currents and the application of the Kubo formula. In this framework, the presence of the axionic field in breaking translational symmetry is crucial for modeling realistic condensed matter systems with momentum dissipation. In our case, the $\sigma_{\tiny{\mbox{DC}}}$ is influenced at the moment to consider the effects of rotation, represented through the parameter $\alpha$, as shown in Fig. 3.
Some interesting directions for future research include, for example, the extension of spinning configurations following this toy model in higher dimensions and exploring a broader range of values for the non-minimal coupling constant $\xi$. Additionally, we can obtain a modification of the first law of BH thermodynamics, considering now the cosmological constant $\Lambda$ as a thermodynamic parameter related to the pressure $P$ of the system Kastor:2009wy ; Kubiznak:2014zwa ; Kubiznak:2016qmn ; Dolan:2010ha ; Cvetic:2010jb , where now the first law takes the form
@@
Here, $V$ corresponds to the conjugate thermodynamic volume and $\cal{M}$ now takes the role of the enthalpy.

### v2

@@ IV Holographic superconductor
In this section, we are interested in exploring a holographic superconductor description of the preceding novel rotating axionic AdS4 BH, investigating how holographic superconductivity behaves in the presence of both rotation and momentum dissipation. To our knowledge, this particular combination of rotation, momentum dissipation, and non-minimal scalar coupling has not been explored before. Given this, we find it convenient to focus on a particular case of this background in which the transformation $u\to r_{h}/r$ makes it feasible to locate the outer horizon at $u=1$ and the asymptotic boundary at $u=0$. This is possible, provided we define
@@
which preserves the genuine nature of the constant parameter $\alpha$ as a rotation parameter. In this manner, by evaluating the above expressions into the background functions (19)-(23) and taking the limit $\alpha\to 0^{+}$, the angular momentum of the background is turned off and a static AdS4 BH configuration is obtained (see eqs. (24)-(25)). Moreover, in the setting (52), the Hawking temperature (49) adopts the simpler form
@@
in which $F_{\mu\nu}=2\partial_{[\mu}A_{\nu]}$, while $\Psi$ denotes a complex scalar field characterized by mass $m$ and charge $q$, minimally coupled to the gauge field $A_{\mu}$ through the extended covariant derivative $D_{\mu}=\nabla_{\mu}-iA_{\mu}$. Here, we work in the probe limit $q\gg 1$, where the backreaction of the matter contribution (54) onto the background fields can be effectively ignored. Thus, the coupled system of field equations is given by
@@
which is intended to be solved on the fixed background (6), taking into account eq. (52). This coupling system is a direct consequence of the structure of the background solution, where the scalar and axionic sectors interact with the metric. Once the background is fixed, the system is solved simultaneously with the standard boundary conditions at the horizon and at the AdS boundary.
@@ IV.1 Condensation of the scalar field and the electric conductivity
where now the primes $(^{\prime})$ stand for derivatives with respect to $u$. Eqs. (57a)-(57c) remain coupled because the function $N^{\theta}$ mixes temporal and spatial components of the gauge field $A_{\mu}$. This is a characteristic feature of rotating holographic backgrounds, and it is precisely this mixing that allows the parameter $\alpha$ to influence the superconducting phase. In solving this system of equations, we need to specify appropriate boundary conditions at the horizon ($u=1$) and at the conformal boundary of the background ($u=0$). Concretely, at the horizon, we require that the scalar field is finite $\Psi(1)<\infty$, while the components $A_{t}(1)$ and $A_{\theta}(1)$ vanish for $A_{\mu}$ to have a finite norm everywhere. On the other hand, at the boundary, the fields behave as
@@
From the holographic standpoint, the coefficient $\Psi_{1}$ (respectively $\Psi_{2}$) is identified as the order parameter $\mathcal{O}$ with scaling dimension $\Delta_{1}$ (respectively $\Delta_{2}$). However, to have well-defined normalizable modes at the Ultraviolet (UV) regime (boundary), we must restrict the scalar field mass to be bounded below as $m^{2}\ell^{2}\geq-\frac{9}{4}$, which defines the Breitenlohner-Freedman (BF) bound for a massive scalar field in AdS4 Breitenlohner:1982bm ; Breitenlohner:1982jf . Besides, $\mu$ and $\rho$ are respectively interpreted as the chemical potential and the charge density operator. A similar interpretation is given to the coefficients of $A_{\theta}$, where $\mu_{\theta}$ stands for a one-dimensional potential vector, the source of the density of the current operator $\mathcal{J_{\theta}}$.
@@
Using the holographic dictionary and by numerically solving the system (57), we determine the condensation profile of the order parameter as a function of the temperature. For our analysis, we work in the canonical ensemble, keeping the charge density $\rho$ fixed. Moreover, we set the scalar mass to $m^{2}\ell^{2}=-2$, a value consistent with the BF bound, and that ensures the simplest asymptotic structure for the scalar field. Concretely, this choice sets the scaling dimensions of the dual operators to $\Delta_{1}=1$ and $\Delta_{2}=2$ for $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$, respectively.
@@
Figura 2: The condensation profiles for the operators $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ as a function of temperature, considering different values of the rotation parameter $\alpha$. The continuous (black) curve corresponds to the static case, while the dashed (violet), dot-dashed (blue), and dotted (red) curves correspond to situations with non-zero rotation.
In Fig. 2, we illustrate the influence of the background rotation on the condensation curves for $\mathcal{O}_{1}$ (left panel) and $\mathcal{O}_{2}$ (right panel). We observe that, as the BH rotation increases, the amplitudes of the condensation profiles decrease. This suppression suggests that rotational effects act as an effective mechanism opposing the formation of the superconducting phase. Such behavior is consistent with previous holographic studies employing rotating BH backgrounds, though in asymptotically different spacetimes, which have likewise reported a reduction of the condensate with increasing rotation Lin:2014tza ; Herrera-Mendoza:2024vfj ; Srivastav:2019ixc . Our results reinforce the robustness of this phenomenon across distinct holographic setups, highlighting the role of the background rotation in the condensate.
In addition to the condensation analysis, we now examine the electric conductivity $\sigma_{\tiny{\mathrm{DC}}}$, which encodes the linear response of the dual system to small electric perturbations. In the holographic superconductor framework of Refs. Hartnoll:2008vx ; Hartnoll:2008kx , the optical conductivity is obtained by turning on time-dependent perturbations of the bulk gauge field and extracting the retarded Green’s function of the boundary current. In our setup, we consider fluctuations of the spatial components $A_{\theta}$ and $A_{\varphi}$, denoted by $\delta A_{\theta}$ and $\delta A_{\varphi}$, at the AdS boundary. Using the ansätze (see Refs. Melnikov:2012tb ; Hartnoll:2008vx )
$\delta A_{\theta,\varphi}=A_{\theta,\varphi}(r)e^{-i\omega t},\qquad\delta\Psi=0,$
where $A_{\theta,\varphi}$ jointly denotes the $\theta$ and $\varphi$ components and $\omega$ is the frequency, and introducing the complex fluctuation $A=A_{\theta}d\theta+iA_{\varphi}d\varphi$, the Maxwell equations $\nabla_{\mu}F^{\mu\nu}=0$ reduce on the rotating axionic background (6) to:
$A^{\prime\prime}_{\theta,\varphi}+\frac{1}{2}\left(\frac{f^{\prime}}{f}+\frac{h^{\prime}}{h}\right)A^{\prime}_{\theta,\varphi}+\frac{\omega^{2}}{f\,h}A_{\theta,\varphi}=0,$
@@
The function $h$ encodes the effect of the rotation function $N^{\theta}$, which is supported by the non-minimally coupled scalar field $\phi$ and the axionic field $\psi$. In contrast with the planar Schwarzschild-AdS4 BH of Refs. Hartnoll:2008vx ; Hartnoll:2008kx , here the rotation and axion sector introduce a non-trivial mixing of temporal and spatial directions and provide an intrinsic mechanism for momentum dissipation in the boundary theory. Following the standard linear-response prescription Hartnoll:2008vx ; Hartnoll:2008kx , an external oscillating electric field $E_{i}(\omega)$ (with $i=\{\theta,\varphi\}$) induces a boundary current $J_{i}(\omega)$, and the complex conductivity is defined as
$\sigma_{\tiny{\mbox{DC}}}(\omega)=\frac{J_{\theta,\varphi}(\omega)}{E_{\theta,\varphi}(\omega)}.$
Holographically, the electric conductivity is related to the retarded Green’s function of the boundary current via the Kubo formula Baggioli:2016rdj :
@@
$\sigma_{\tiny{\mbox{DC}}}(\omega)=-\frac{i}{\omega}G^{R}_{J_{\theta,\varphi}J_{\theta,\varphi}}(\omega)=\frac{A^{\prime}_{\theta,\varphi}}{i\omega A_{\theta,\varphi}}\Big|_{r\to+\infty}.$
Here, the last equality follows from the standard near-boundary expansion of the bulk gauge field. The ingoing boundary condition at the horizon encodes dissipation in the dual field theory. In contrast, the asymptotic expansion at the AdS boundary determines the source (electric field) and the response (current).
@@
Figura 3: The $\sigma_{\tiny{\mbox{DC}}}$ curve for the real (left panel) and imaginary (right panel) parts of the electrical conductivity for different values of the rotation parameter $\alpha$. The continuous (black) curve corresponds to the static case, while the other values correspond to situations with non-zero angular momentum.
Figure 3 shows the real and imaginary parts of $\sigma_{\tiny{\mbox{DC}}}$ for different values of the rotation parameter $\alpha$. For $\alpha=0$ (the continuous black curve), the behavior is qualitatively similar to that of the original planar holographic superconductor (see Refs. Hartnoll:2008vx ; Hartnoll:2008kx ; Gubser:2008px ; Horowitz:2008bn ) where the real part develops a pronounced low-frequency enhancement, while the imaginary part displays a simple pole at $\omega=0$. For the situation $\alpha\neq 0$, the rotation deforms this pattern. The pole in $\mathrm{Im}(\sigma_{\tiny{\mbox{DC}}}(\omega))$ at $\omega=0$ persists, signalling a $\delta(\omega)$ contribution in $\mathrm{Re}(\sigma_{\tiny{\mbox{DC}}})$ through the Kramers–Kronig relations and thus the presence of a dissipationless component of the current, associated with the holographic superconducting phase, exactly as in Refs. Hartnoll:2008vx ; Hartnoll:2008kx ; Hartnoll:2009sz . In the present model, however, the origin of this pole is not solely tied to momentum conservation. The axionic sector breaks translational invariance, leading to momentum relaxation in the dual theory and allowing charge carriers to dissipate their momentum into the effective “lattice” generated by the axion profile, in close analogy with linear-axion holographic models of momentum dissipation Andrade:2013gsa . From a condensed-matter perspective, this mimics the effects of impurities, crystalline lattices, and disorder Amoretti:2019cef ; Ammon:2019wci ; Andrade:2017lsc . The rotating axionic background thus provides a controlled holographic realization of a superconductor with both rotation and momentum dissipation. As the constant parameter $\alpha$ increases, the low-frequency structure of $\mathrm{Re}(\sigma_{\tiny{\mbox{DC}}})$ and the strength of the $1/\omega$ pole in $\mathrm{Im}(\sigma_{\tiny{\mbox{DC}}})$ are modified, reflecting a non-trivial interplay between rotation, axion-induced momentum relaxation, and the superconducting condensate. This qualitative behavior is consistent with the suppression of the condensate by rotation observed in other holographic models based on rotating BHs Herrera-Mendoza:2024vfj ; Lin:2014tza ; Srivastav:2019ixc , but arises here in a background with non-minimal scalar coupling and an explicit axionic momentum-dissipation mechanism.
Therefore, the numerical profiles in Fig. 3 show that the rotating axionic BH provides a genuinely new setting for holographic superconductivity. The conductivity encodes how the rotation constant parameter $\alpha$, the axionic sector and the scalar couplings jointly determine the transport properties of the dual strongly coupled system, extending the original planar Schwarzschild–AdS4 model of Hartnoll:2008vx ; Hartnoll:2008kx ; Gubser:2008px ; Horowitz:2008bn to a rotating, momentum-relaxing background, in the broader spirit of holographic transport and strange-metal phenomenology.
@@ V Conclusions and discussions
In this work, we have constructed a new axionically charged spinning BH in four dimensions with a non-minimally coupled scalar field. The above is possible via the inclusion of a suitable potential as well as a structural function coupling to the axionic field. Although the metric function $f$ and the rotating function $N^{\theta}$ enjoy logarithmic behavior, this configuration is asymptotically AdS. These configurations extend previous three-dimensional results and open avenues for further analysis of thermodynamics, stability, and holographic applications.
As can be seen in Fig. 1 (left panel), this configuration is characterized by two constant parameters ($\alpha$ and $\eta$) and only one positive integration constant $B$, allowing us to obtain one or even two horizons. Furthermore, we find that the metric function $f$ and the scalar field $\phi$ do not vanish for $\alpha\rightarrow 0^{+}$ and $\eta\neq 0$. This allows us to obtain a new static hairy BH configuration with one horizon (see Fig. 1 (right panel)).
The presence of the scalar and axionic fields introduces additional thermodynamic variables. Using the Euclidean action (34), we derived the mass $\mathcal{M}$, Hawking temperature $T$, angular momentum $\mathcal{J}$, angular velocity $\Omega$, as well as the axionic potential $\Psi_{a}$, and axionic charge $\omega_{a}$, satisfying the first law (51). It is straightforward to check that upon the following dimensional scaling $r_{h}\rightarrow\tilde{\lambda}r_{h}$ and $B\rightarrow\tilde{\lambda}B$, we have:
${\cal{M}}\rightarrow\tilde{\lambda}^{3}{\cal{M}},\quad{\cal{S}}\rightarrow\tilde{\lambda}^{2}{\cal{S}},\quad{\cal{J}}\rightarrow\tilde{\lambda}^{3}{\cal{J}},\quad\omega_{a}\rightarrow\tilde{\lambda}\omega_{a},$
yielding the four-dimensional Smarr relation Smarr:1972kt
@@
In exploring the holographic superconductor description of these spinning BHs, we employed a numerical approach to investigate the condensation of the order parameter as a function of temperature and the rotation parameter (see Fig. 2). Notably, it is observed that the presence of a logarithmic falloff in the blackening function does not influence the characteristic behavior of the condensate. Nevertheless, our results reveal that the BH rotation impacts the order parameter. Specifically, we found that an increase in the BH rotation corresponds to a suppression in the amplitudes of the condensation curves, suggesting that the rotation is responsible for inhibiting the superconducting phase in our holographic system. This effect becomes relevant only at sufficiently high values of the rotation parameter.
Together with the aforementioned, the electric conductivity as a function of the frequency $\omega$ is calculated numerically from the spinning configuration through the response of the system to small external sources. This is followed by the calculation of the Green’s functions for the electric currents and the application of the Kubo formula. In this framework, the presence of the axionic field in breaking translational symmetry is crucial for modeling realistic condensed matter systems with momentum dissipation. In our case, $\sigma_{\tiny{\mbox{DC}}}$ is influenced at the moment to consider the effects of rotation, represented through the parameter $\alpha$, as shown in Fig. 3.
Some interesting directions for future research include, for example, the extension of spinning configurations following this toy model in higher dimensions and exploring a broader range of values for the non-minimal coupling constant $\xi$. Axion BHs solutions have also been constructed in the context of effective string theory, most notably in the works of Gal’tsov, Kechkin et al. Galtsov:1994pd ; Galtsov:1995zu ; Garcia:1995qz . Although these configurations allow us valuable insights into the role of axionic fields within gravitational backgrounds, their construction is based on the absence of a cosmological constant and relies on specific solution-generating symmetries that do not directly apply to the AdS case considered in the present work. Establishing corresponding solutions in the AdS context would necessitate a separate analysis. Nonetheless, it would be intriguing to explore these generalizations in future research. Additionally, we can obtain a modification of the first law of BH thermodynamics, considering now the cosmological constant $\Lambda$ as a thermodynamic parameter related to the pressure $P$ of the system Kastor:2009wy ; Kubiznak:2014zwa ; Kubiznak:2016qmn ; Dolan:2010ha ; Cvetic:2010jb , where now the first law takes the form
@@
[…]